Can you help me solve the equation for x
5-3(2x-7)=8-4x

Answer by pradhyumna.kumar@gmail.com(62)

Can you help me solve the equation for x
5-3(2x-7)=8-4x
sol : : 5 - 6x + 21 = 8 - 4x
18 = 2x => x = 9